OpenStudy (anonymous):

3×√2−√5/3+√2−√5=a+b×√5 then find a-b

OpenStudy (anonymous):

\[\frac{ 3\times \sqrt{2}-\sqrt{5} }{ 3+\sqrt{2}-\sqrt{5} }=a+b \times \sqrt{5}\] then find a-b

OpenStudy (anonymous):

hey some one help

OpenStudy (anonymous):

Are you sure the question you are given is true?

OpenStudy (anonymous):

yaa but its 3into(x) not x

OpenStudy (anonymous):

Is it this?